# On random matrix statistics of 3d gravity

### v1

@@ Abstract
We show that 3d gravity on manifolds that are topologically a Riemann surface times an interval $\Sigma_{g,n}\times I$ with end-of-the-world branes at the ends of the interval is described by a random matrix model, namely the Virasoro minimal string. Because these manifolds have $n$ annular asymptotic boundaries, the path integrals naturally correspond to spectral correlators of open strings upon inverse Fourier transforms. For $g=0$ and $n=2$, we carry out an explicit path integration and find precise agreement with the universal random matrix expression. For Riemann surfaces with negative Euler characteristic, we evaluate the path integral as a gravitational inner product between states prepared by two copies of Virasoro TQFT. Along the way, we clarify the effects of gauging the mapping class group and the connection to chiral 3d gravity.
@@ 1 Introduction
The gravitational path integral is notoriously difficult to evaluate in general. In three-dimensional pure Einstein gravity with a negative cosmological constant (3d gravity), however, substantial progress has been made. For manifolds admitting a hyperbolic structure, the formulation in terms of Virasoro topological quantum field theory (Virasoro TQFT) Collier:2023fwi ; Collier:2024mgv provides a conceptually simple method for computing the path integral at finite gravitational coupling. Such manifolds have played an important role in revealing the ensemble nature of CFTs dual to 3d gravity Kraus:2016nwo ; Cardy:2017qhl ; Collier:2019weq ; Belin:2020hea ; Belin:2021ryy ; Anous:2021caj ; Kusuki:2021gpt ; Numasawa:2022cni ; Chandra:2022bqq ; deBoer:2023vsm ; deBoer:2024mqg ; Wang:2025bcx ; Hung:2025vgs ; Jafferis:2025yxt .
@@
Classically, 3d gravity is equivalent to a Chern-Simons theory Achucarro:1986uwr ; Witten:1988hc . This is expected to extend to the quantum theory perturbatively Witten:2007kt ; Kim:2015qoa . Non-perturbatively, the two theories are different. Nevertheless, by requiring the gauge fields to correspond to non-degenerate metrics, one picks out the Teichmüller component Krasnov:2005dm ; Scarinci:2011np . With this restriction, the Chern-Simons theory is equivalent to 3d gravity, except that large diffeomorphisms must be gauged in gravity. For hyperbolic manifolds, the effect of gauging the mapping class group is rather trivial, but the same cannot be said about off-shell manifolds. For example, Virasoro TQFT gives a divergent path integral for the torus wormhole Collier:2023fwi , while the direct computation gives a finite answer Cotler:2020ugk . It is believed that gauging the mapping class group will resolve this discrepancy Jafferis:2024jkb , though it has not been demonstrated explicitly. In this work, we will see that the annulus wormhole is a simpler setup where we can see the effect of gauging the mapping class group explicitly, turning a divergent expression finite.
@@ 2 Duality via BCFT
where $\mathcal{H}^{(ab)}$ is the Hilbert space of the open-string states, $H^{(ab)}$ is the Hamiltonian generating the direction along the circle, and $2\pi\,{\rm Im}(\tau)=\beta$ is the inverse temperature. The geometric result of taking the trace is an annulus, generally with different boundary conditions on the two circles, $a$ and $b$. The Hilbert spaces with different pairs of boundary conditions are independent.
@@
where $B^{a}$ and $B^{b}$ are boundary states at the circular borders, with $a$ and $b$ labeling the boundary conditions, and $\tilde{H}$ is the Hamiltonian generating the translation in the interval direction. This is often called the closed-string channel.
@@
$\tau=\mathrm{i}t=\mathrm{i}\beta/2\pi.$ (2.6)
@@ 3 Disk amplitude
which is the vacuum character. We then use (2.5) to write the open-channel expression as
@@ 4 Cylinder amplitude
Noticing that $A_{y}$ and $\bar{A}_{y}$ appear as Lagrange multipliers in the action, we path integrate them first, setting the spatial field strength $F_{x\rho}=\bar{F}_{x\rho}=0$. The flatness conditions means that the connections can be written as
@@
with $G$ single-valued. In terms of $G$,
@@
Now, let us explain the key observation that will simplify the calculation. In analogy with the CFT doubling trick of Cardy Cardy:1984bb , which has a natural uplift to Chern-Simons theory, we perform the following actions, as illustrated by the diagrams (showing only the $y$ direction):
@@
The non-trivial aspect of this redefinition has to do with the EOW boundary conditions and the action. By (4.4), the field $A$ is continuous across $y=0,\pi$, so it remains a valid configuration in the doubled manifold. As for the action, notice that the $\bar{A}$ part of the bulk action in (4.11) flips sign under this redefinition, while the boundary action remains invariant. So we can simply replace the action with
@@
where we have used the fact that $\tau_{i}$ is purely imaginary for us, and the labels 1 and 2 now refer to the two asymptotic boundaries of the doubled manifold (each of topology $A\times S^{1}$). This has turned the problem into one looking like the torus case, but with only a chiral half of the action.
@@
where we have dropped the label under the integral sign here and in what follows, and it should be understood that all $y$ integrals will range between 0 and $2\pi$ unless otherwise stated.
As $x$ is a compact direction, the total $x$-derivative terms would vanish for periodic fields. (The only field non-periodic in $x$ here is $\Phi=\phi+b(y)x$.) This expression therefore simplifies significantly, in fact reducing to a purely boundary term (presenting only the contribution at boundary 1)
@@
The boundary action is (presenting only for boundary 1)
@@
$f_{n}(-\infty)=0,\quad f_{n}(\infty)=2\pi n,\quad n\in\mathbb{Z}.$ (4.42)
@@
where $b^{2}_{n}$ are the Fourier coefficients of $b^{2}(y)$ rather than the squares of the Fourier coefficients of $b(y)$.
@@
$\langle\tilde{Z}(\tau_{1})\tilde{Z}(\tau_{2})\rangle=2\pi\mathcal{N}\int\mathrm{d}b^{2}Z_{\rm AS}(\tau_{1}|b)Z_{\rm AS}(\tau_{2}|b).$ (4.64)
@@
$\mathcal{Z}_{A\times I}(\mathrm{i}t_{1},\mathrm{i}t_{2})$ $=\langle{Z}(\mathrm{i}t_{1}){Z}(\mathrm{i}t_{2})\rangle=\langle\tilde{Z}(\mathrm{i}/t_{1})\tilde{Z}(\mathrm{i}/t_{1})\rangle$
@@
We have implicitly assumed that $a$ and $b$ are different boundary conditions, so there is only one way to connect the two asymptotic annuli. In the special case $a=b$, i.e., with identical boundary conditions, there are two ways to connect the annuli, so we get a factor of two, making it a GOE ensemble Bohigas:1983er . This is in fact the correct answer, as the matrix potential is a sum over $a$ and $b$, so the diagonal terms have a factor of a half relative to off-diagonal terms Jafferis:2025yxt , making it GOE when $a=b$.
@@ 5 Other amplitudes
The partition function of 3d gravity on $M=\Sigma_{g,n}\times I$, with asymptotic boundary conditions at the ends of the interval, is equal to the Liouville CFT partition function on $\Sigma_{g,n}$ Chandra:2022bqq ; Collier:2023fwi . It is sometimes convenient to think of the Liouville partition function as a state on $\Sigma_{g,n}$ with moduli $(\mathbf{m}_{1},{\mathbf{m}_{2}})$ prepared by two copies of Virasoro TQFT on the $\mathbb{Z}_{2}$ quotient of the Maldacena-Maoz (MM) wormhole Jafferis:2025yxt :
@@
It is useful to rewrite the phase space $(\mathcal{T}_{g,n}\times\mathcal{T}_{g,n})/\text{Map}(\Sigma_{g,n})$ as a $\mathcal{T}_{g,n}$ non-trivial bundle over $\mathcal{M}_{g,n}$, where we use ${\mathbf{m}}_{2}$ for coordinates on the fiber and ${\mathbf{m}}_{1}$ for coordinates on the base. At each point in the base, we then perform the integration along the fiber:
@@
where $\mathsf{Z}_{g,n}\left(\beta_{1},\ldots,\beta_{n}\right)$ is the VMS partition function on $\Sigma_{g,n}$ with (circular) asymptotic boundaries with inverse temperatures $\beta_{i}$, with the $\mathrm{e}^{S_{0}}$ dependence kept.
@@ 6 Discussion
We have restricted our attention to a large class of off-shell manifolds with EOW branes, namely $\Sigma_{g,n}\times I$, and shown that they are related to a random matrix model. Other off-shell manifolds can be obtained by starting with on-shell manifolds and gluing these manifolds to them Jafferis:2024jkb . In the purely open tensor-matrix model of Jafferis:2025yxt , on-shell manifolds with EOW branes and boundary Wilson loops are generated by the tensor part of the model’s potential. We then remove the neighborhood of each of the Wilson loops, leaving an annular state-cut boundary. At the leading order, the thickened disk (3.1) is glued to each of such annuli, a procedure also known as the annular surgery. For any set of $n$ annuli, we glue $\Sigma_{g,n}\times I$ to it for any $g$ to obtain higher-order corrections. This makes predictions for the partition functions of manifolds with more complicated off-shell topology, and it would be interesting to check them.
@@
Finally, it is noteworthy that when the asymptotic boundary is a single annulus rather than a torus, there is no $\mathrm{SL}(2,\mathbb{Z})$ modular sum, owing to the fact that the annulus does not have such symmetries. The spectral density is manifestly positive and does not suffer from the negativity issue encountered in Maloney:2007ud ; Keller:2014xba . This hopefully provides some hints on how to fix the negativity issue for the closed-string sector.
@@ Acknowledgements
It is a pleasure to thank Scott Collier, Tom Hartman, Alex Maloney, Julian Sonner, Joaquin Turiaci, Zixia Wei, Cynthia Yan, and Mengyang Zhang for helpful discussions. DLJ, LR and DS acknowledge support by the Simons Investigator in Physics Award MP-SIP-0001737, U.S. Department of Energy grant DE-SC0007870, and Harvard University. DW acknowledges support by NSF grant PHY-2207659 and the Simons Collaboration on Celestial Holography.

### v2

@@ Abstract
We show that 3d gravity on manifolds that are topologically a Riemann surface times an interval $\Sigma_{g,n}\times I$ with end-of-the-world branes at the ends of the interval is described by a random matrix model, namely the Virasoro minimal string. Because these manifolds have $n$ annular asymptotic boundaries, the path integrals naturally correspond to spectral correlators of open strings upon inverse Laplace transforms. For $g=0$ and $n=2$, we carry out an explicit path integration and find precise agreement with the universal random matrix expression. For Riemann surfaces with negative Euler characteristic, we evaluate the path integral as a gravitational inner product between states prepared by two copies of Virasoro TQFT. Along the way, we clarify the effects of gauging the mapping class group and the connection to chiral 3d gravity.
@@ 1 Introduction
The gravitational path integral is notoriously difficult to evaluate in general. In three-dimensional pure Einstein gravity with a negative cosmological constant (3d gravity), however, substantial progress has been made. For manifolds admitting a hyperbolic structure, the formulation in terms of Virasoro topological quantum field theory (Virasoro TQFT) Collier:2023fwi ; Collier:2024mgv provides a conceptually simple method for computing the path integral at finite gravitational coupling. Such manifolds have played an important role in revealing the ensemble nature of CFTs dual to 3d gravity Kraus:2016nwo ; Cardy:2017qhl ; Collier:2019weq ; Belin:2020hea ; Belin:2021ryy ; Anous:2021caj ; Kusuki:2021gpt ; Schlenker:2022dyo ; Numasawa:2022cni ; Chandra:2022bqq ; deBoer:2023vsm ; deBoer:2024mqg ; Wang:2025bcx ; Hung:2025vgs ; Jafferis:2025yxt ; Chandra:2025fef ; Hartman:2025ula ; Liu:2025ikq .
@@
Classically, 3d gravity is equivalent to a Chern-Simons theory Achucarro:1986uwr ; Witten:1988hc . This is expected to extend to the quantum theory perturbatively Witten:2007kt ; Kim:2015qoa . Non-perturbatively, the two theories are different. Nevertheless, by requiring the gauge fields to correspond to non-degenerate metrics, one picks out the Teichmüller component Krasnov:2005dm ; Scarinci:2011np , and with this restriction, the Chern-Simons theory is equivalent to 3d gravity, except that large diffeomorphisms must be gauged in gravity. For hyperbolic manifolds, the effect of gauging the mapping class group is rather trivial, but the same cannot be said about off-shell manifolds. For example, Virasoro TQFT gives a divergent path integral for the torus wormhole Collier:2023fwi , while the direct computation gives a finite answer Cotler:2020ugk . It is believed that gauging the mapping class group will resolve this discrepancy Jafferis:2024jkb , though it has not been demonstrated explicitly. In this work, we will see that the annulus wormhole is a simple setup where we can see the effect of gauging the mapping class group explicitly, rendering a divergent expression finite.
@@ 2 Duality via BCFT
where $\mathcal{H}^{(ab)}$ is the Hilbert space of the open-string states, $H^{(ab)}$ is the Hamiltonian generating the direction along the circle, and $2\pi\,{\rm Im}(\tau)=\beta$ is the inverse temperature. The geometric result of taking the trace is an annulus, generally with different boundary conditions on the two circular borders, $a$ and $b$. The Hilbert spaces with different pairs of boundary conditions are independent.
@@
where $B^{a}$ and $B^{b}$ are boundary states at the circular borders, with $a$ and $b$ labeling the boundary conditions, and $\tilde{H}$ is the Hamiltonian generating translation in the interval direction. This is often called the closed-string channel.
@@
$\tau\equiv\mathrm{i}t=\mathrm{i}\beta/2\pi.$ (2.6)
@@ 3 Disk amplitude
which is the vacuum character multiplied by $g$-functions. We then use (2.5) to write the open-channel expression as
@@ 4 Cylinder amplitude
Noticing that $A_{y}$ and $\bar{A}_{y}$ appear as Lagrange multipliers in the action, we path integrate them first, setting the spatial field strength $F_{x\rho}=\bar{F}_{x\rho}=0$.
The flatness conditions means that the connections can be written as
@@
with $G$ and $\bar{G}$ single-valued. In terms of $G$,
@@
Now, let us explain the key fact that will simplify the calculation. In analogy with the CFT doubling trick of Cardy Cardy:1984bb , which has a natural uplift to Chern-Simons theory, we perform the following actions, as illustrated by the diagrams (showing only the $y$ direction):
@@
The non-trivial aspect of this redefinition has to do with the EOW boundary conditions and the action. By (4.4), the field $A$ is continuous across $y=0,\pi$, so it remains a valid configuration in the doubled manifold. As for the action, notice that the $\bar{A}$ part of the bulk action in (4.11) flips sign under this redefinition, while the boundary action remains invariant. So we can simply replace the action (4.11) with
@@
where we have used the fact that $\tau_{i}$ is purely imaginary for us, and the labels 1 and 2 now refer to the two asymptotic boundaries of the doubled manifold (each of topology $I\times S^{1}$). This has turned the problem into one looking like the torus case, but with only a chiral half of the action. The problem should now be more familiar to us, as this is just the action of chiral 3d gravity Li:2008dq . We are free to confuse the two theories as far as the action is concerned, though as we shall see, it is important that we are working with 3d gravity with EOW branes, not chiral 3d gravity, when discussing global issues such as the mapping class group.
@@
where we have dropped the label under the integral sign here, and it should henceforth be understood that all $y$ integrals will range between 0 and $2\pi$ unless otherwise stated.
As $x$ is a compact direction, the total $x$-derivative terms would vanish for periodic fields. (The only field non-periodic in $x$ here is $\Phi=\phi+b(y)x$.) This expression therefore simplifies significantly, in fact reducing to a purely boundary term (presenting only the contribution at boundary 1):
@@
The boundary action is (again presenting only for boundary 1)
@@
where $f_{n}(-\infty)=0,\quad f_{n}(\infty)=2\pi n,\quad n\in\mathbb{Z}.$ (4.42)
@@
where $b^{2}_{n}$ are the Fourier coefficients of $b^{2}(y)$ (not the squares of the Fourier coefficients of $b(y)$, despite what the notation might suggest).
@@
$\langle\tilde{Z}(\tau_{1})\tilde{Z}(\tau_{2})\rangle_{0}=2\pi\mathcal{N}\int\mathrm{d}b^{2}Z_{\rm AS}(\tau_{1}|b)Z_{\rm AS}(\tau_{2}|b).$ (4.64)
@@
$\mathcal{Z}_{A\times I}(\mathrm{i}t_{1},\mathrm{i}t_{2})$ $=\langle{Z}(\mathrm{i}t_{1}){Z}(\mathrm{i}t_{2})\rangle_{0}=\langle\tilde{Z}(\mathrm{i}/t_{1})\tilde{Z}(\mathrm{i}/t_{1})\rangle_{0}$
@@
We have implicitly assumed that $a$ and $b$ are different boundary conditions, so there is only one way to connect the two asymptotic annuli. However, in the special case $a=b$, i.e., with identical boundary conditions, there are two ways to connect the annuli. To be more precise, the BCFT annulus partition has a CRT symmetry which exchanges the two circular borders. Similar to the closed case Yan:2023rjh , this means that we can cut the wormhole in the middle and glue them back together after the $\mathbb{Z}_{2}$ mapping group class action swapping the two EOW branes:
[…]
This give a factor of two for the case $a=b$ relative to when $a\neq b$, making it a GOE ensemble Bohigas:1983er . This is in fact predicted by the matrix potential, which is a sum over $a$ and $b$, so the diagonal terms have a factor of a half relative to off-diagonal terms, which becomes a factor of two after taking the inverse Jafferis:2025yxt .
Before moving on, we shall emphasize a point regarding the mapping class group: when considering the mapping class group, we should fix the EOW boundaries set-wise, as they are dynamical objects in the theory. This is in contrast to asymptotic boundaries, which we fix point-wise. In (4), the mapping class group of the annulus at $y=0$ is $\mathbb{Z}$ because we fix the asymptotic boundaries point-wise. When considering the annulus at $\rho=0$ as in (4.68), the mapping class group of the annulus is trivial (when $a\neq b$) rather than $\mathbb{Z}$ because the EOW circles are fixed set-wise.
@@ 5 Other amplitudes
The partition function of 3d gravity on Maldacena-Maoz (MM) wormholes Jafferis:2025yxt , which are topologically $M=\Sigma_{g,n}\times I$, with asymptotic boundary conditions at the ends of the interval, is equal to the Liouville CFT partition function on $\Sigma_{g,n}$ Chandra:2022bqq ; Collier:2023fwi . It is sometimes convenient to think of the Liouville partition function as a state on $\Sigma_{g,n}$ with moduli $(\mathbf{m}_{1},{\mathbf{m}_{2}})$ prepared by two copies of Virasoro TQFT on the $\mathbb{Z}_{2}$ quotient of the MM wormhole Jafferis:2025yxt :
@@
It is useful to rewrite the phase space $(\mathcal{T}_{g,n}\times\mathcal{T}_{g,n})/\text{Map}(\Sigma_{g,n})$ as a (non-trivial) $\mathcal{T}_{g,n}$ bundle over $\mathcal{M}_{g,n}$, where we use ${\mathbf{m}}_{2}$ for coordinates on the fiber and ${\mathbf{m}}_{1}$ for coordinates on the base. At each point in the base, we then perform the integration along the fiber:
@@
where $\mathsf{Z}_{g,n}\left(\beta_{1},\ldots,\beta_{n}\right)$ is the VMS partition function on $\Sigma_{g,n}$ with (circular) asymptotic boundaries with inverse temperatures $\beta_{i}$, with the $\mathrm{e}^{S_{0}}$ dependence kept Collier:2023cyw .
@@ 6 Discussion
We have restricted our attention to a large class of off-shell manifolds with EOW branes, namely $\Sigma_{g,n}\times I$, and shown that they are related to a random matrix model. Other off-shell manifolds can be obtained by starting with on-shell manifolds and gluing these manifolds to them Jafferis:2024jkb . In the purely open tensor-matrix model of Jafferis:2025yxt , on-shell manifolds with EOW branes and boundary Wilson loops are generated by the tensor part of the model’s potential. We then remove the neighborhood of each of the Wilson loops, leaving an annular state-cut boundary. At the leading order, the thickened disk (3.1) is glued to each of such annuli, a procedure also known as the annular surgery. For any set of $n$ annuli, we glue $\Sigma_{g,n}\times I$ to it for every $g$ to obtain higher-order corrections. This makes predictions for the partition functions of manifolds with more complicated off-shell topology, and it would be interesting to check them.
@@
Finally, it is noteworthy that when the asymptotic boundary is a single annulus, there is no $\mathrm{SL}(2,\mathbb{Z})$ modular sum like the torus case. A annulus that is homotopic to the asymptotic boundary (say a constant-radius slice) has trivial mapping class group (since it has EOW boundaries, as explained at the end of Section 4). The analog of the Maloney-Witten sum Maloney:2007ud ; Keller:2014xba is therefore just the Cardy density (times the $g$-functions) plus the vacuum state. This spectral density is manifestly positive.
@@ Acknowledgements
It is a pleasure to thank Jeevan Chandra, Scott Collier, Gabriele Di Ubaldo, Tom Hartman, Luca Iliesiu, Alex Maloney, Julian Sonner, Joaquin Turiaci, Zixia Wei, Cynthia Yan, and Mengyang Zhang for helpful discussions. DLJ, LR and DS acknowledge support by the Simons Investigator in Physics Award MP-SIP-0001737, U.S. Department of Energy grant DE-SC0007870, and Harvard University. DW acknowledges support by NSF grant PHY-2207659 and the Simons Collaboration on Celestial Holography.